OpenStudy (anonymous):

A jug will hold 2/3 gallon of punch. How much punch is in the jug when it is 3/4 full?

OpenStudy (anonymous):

\[\frac{3}{4}\left(\frac{2}{3}\right)=\frac{1}{2} \text{gallon} \]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

You're welcome.